Parth (parthkohli):

Do you think that Fermat would have been able to successfully prove his Last Theorem if he had not died?

Parth (parthkohli):

I don't think he would've done it. It was a long proof, after all.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

No

Parth (parthkohli):

I have support. Wow.

OpenStudy (anonymous):

sometimes u can see here http://en.wikipedia.org/wiki/Fermat's_Last_Theorem

Parth (parthkohli):

Please read the question again, @best.shakir

OpenStudy (anonymous):

i read the question but not the answer

mathslover (mathslover):

Yes @ParthKohli He would have .. I think

Parth (parthkohli):

How so?

mathslover (mathslover):

his aim .. he would have like to achieve that?

Parth (parthkohli):

He actually said that the theorem existed just for nothing; he didn't know the proof.

mathslover (mathslover):

Sometime people get ideas at any instant? I mean may be he was not able to do that before but he would have been after

OpenStudy (amistre64):

from what i remember of fermat .... yes

OpenStudy (anonymous):

I believe he thought he had a proof, but it was flawed and he didn't see it.

Parth (parthkohli):

He was a bit of a genius, but he still didn't know the theorem well.

OpenStudy (vishweshshrimali5):

@satellite73 , just want to know if this theorem is really true............ I don't have much about this. so please....

Parth (parthkohli):

You may ask me, lol. Satellite is not the only genius on this site. Yes it's true.

OpenStudy (amistre64):

I tend to think that the path that leads us here took a different route than the path Fermat was on; and that we simply have to explain it using this road, which is alot bumpier than his :)

OpenStudy (vishweshshrimali5):

@ParthKohli I didn't meant that....... \(a^3 + b^3 = (a+b)(a^2 - ab + b^2)\) Thus, \(a^3 + b^3 = a+b\) if \(a^2 + b^2 - ab = 0\) just a thought....... please tell me if I am wrong.

Parth (parthkohli):

Remember that \(a,b,c\) are supposed to be integers.

OpenStudy (vishweshshrimali5):

Oh sorry ............. misunderstoof.

OpenStudy (vishweshshrimali5):

* misunderstood.......

OpenStudy (vishweshshrimali5):

but \(\large 0^n = 0^n + 0^n\) What now ....... ?

Parth (parthkohli):

'no three positive integers'. 0 is not said to be positive, Vishesh. http://en.wikipedia.org/wiki/Fermat's_Last_Theorem

Parth (parthkohli):

Vishwesh*

OpenStudy (vishweshshrimali5):

Ohhh thanks a lot @ParthKohli It really helped me........... :)

Parth (parthkohli):

My pleasure!